Jim has 28 more nickels than dimes. (Hint: Always let x represent the coin you have the least of . ) The value of the coins is 665 cents or $6.65. Find the number of nickels that Jim has. You MUST first write an equation

Answer by ewatrrr(24785)


Hi,
Let x and (x+28) represent the number of dimes and nickels respectively
10x + 5(x+28) = 665 cents
15x = 665 - 5*28
15x = 525
x = 35, the number of dimes. Number of Nickels is 63 (35+28)
CHECKING our Answer
10*35 + 5*63 = 350 + 315 = 665 cents or $6.65